# OmniSpatial: Towards Comprehensive Spatial Reasoning Benchmark for Vision Language Models

### v1

@@ 5.1 Benchmarking Spatial Reasoning
The development of standardized benchmarks [55, 122, 65, 124, 44, 51] has been critical in assessing the performance of spatial reasoning models. Various studies have introduced innovative benchmarking methodologies [104, 15, 17, 11, 102, 89, 96] to advance spatial reasoning evaluation. For instance, Spatial VQA [26] was among the first to incorporate spatial information into vision-language models, enabling fundamental spatial relationship reasoning. SpatialRGPT-Bench [17] further introduced depth-aware reasoning, utilizing large language models to analyze spatial relationships, such as understanding relative directions and distances between specified regions in 3D scenes [104]. SpatialBench [11] categorized spatial reasoning into various hierarchical levels, extending its applicability to robotic manipulation tasks. RoboSpatial [102] proposes a large-scale template-based spatial relationship dataset, focusing on positional relationships from different perspectives. VSI-Bench [121] combined video data [14, 65, 33, 56, 32] with cognitive maps [72, 4] to simulate human-like spatial cognition and optimize reasoning in dynamic environments. Recently, SoFar [89] introduced the 6-DoF SpatialBench to evaluate the understanding of orientation.
While these benchmarks have made significant contributions, a unified framework encompassing a wide range of complex spatial reasoning tasks remains lacking. Inspired by prior spatial reasoning research [119, 114, 58, 76], we identified several limitations in existing benchmarks, such as reliance on generated images [104, 34, 48, 61, 92, 101], LLM-generated templates [59, 26, 91, 93], and domain-specific focus [117, 19, 21, 12, 28]. These issues hinder their comprehensiveness and real-world applicability. To address these gaps, our study proposes a comprehensive and integrative spatial reasoning benchmark.
@@ 6 Conclusion
In this paper, we introduce OmniSpatial, a benchmark for comprehensive visual–spatial reasoning. OmniSpatial distills spatial cognition into four primary categories—dynamic reasoning, complex spatial logic, spatial interaction, and perspective-taking—spanning 50 fine-grained subtasks and 1.5 K manually-curated question–answer pairs. Extensive experiments show that state-of-the-art proprietary and open-source VLMs peak at 57% accuracy—over 30 points below human performance—struggling especially with geometric reasoning and non-egocentric perspective taking. By exposing these systematic gaps, OmniSpatial provides a rigorous yardstick and a rich error analysis playground for future work on physics- and viewpoint-aware multimodal models.
@@ B.1 Spatial Vision-Language Models
Spatial vision-language models integrate computer vision [1, 37, 80, 90] and natural language processing [6, 115, 9, 2, 107, 108, 128] to enhance machine understanding of spatial relationships. Recent research [88, 15, 17, 11, 102, 123, 89] has increasingly focused on extending vision-language models to support spatial reasoning in dynamic and 3D environments. One of the pioneering works in this domain, SpatialVLM [15], has significantly advanced spatial reasoning by constructing an RGB-D based visual question-answering dataset. These models [45, 45, 97, 115] effectively process multimodal data containing spatial information, serving as a bridge between visual perception and linguistic reasoning.
Further advancements [62, 60, 94, 105, 120, 99, 116, 103] include SpatialRGPT [17], which extends RGB-D spatial understanding by incorporating 3D scene graphs and spatial data to improve inference capabilities. Similarly, SpatialBot [11] explores hierarchical deep reasoning mechanisms to handle depth and spatial structures in complex environments. Recently, SoFar [89] proposed semantic orientation and trained an open-world orientation model to enhance the orientation understanding of VLMs, significantly improving spatial understanding and robotic operation capabilities [18, 25, 73, 131, 113, 43, 74, 27, 42, 89]. However, existing research remains limited in addressing the full complexity and comprehensiveness of spatial reasoning tasks. To bridge this gap, our study aims to develop a more comprehensive benchmark that rigorously evaluates spatial reasoning capabilities.
@@ B.2 Spatial Reasoning in Psychology
In psychology, spatial reasoning refers to an individual’s ability to acquire, organize, utilize, and adapt spatial knowledge, recognized as one of the nine primary reasonings [35]. To systematically characterize this ability, Buckley et al. [10]proposed a factor analysis-based framework, distinguishing between spatial visualization, spatial relations (such as mental rotation), and spatial orientation. Malanchini et al. [69] further identified a strong correlation between spatial orientation and object manipulation skills. Additionally, Newcombe & Shipley [75]introduced a classification system for spatial thinking, dividing spatial reasoning into two dimensions: intrinsic-extrinsic and static-dynamic. Furthermore, many studies also have contributed to the development of frameworks [112, 52, 69, 39, 70] and evaluation methods [29, 111, 110, 40] for spatial reasoning. Their framework highlights the distinction between object-centric spatial properties and external reference frames, offering valuable insights for applications such as navigation and path planning. These psychological frameworks provide complementary perspectives for understanding and assessing spatial reasoning, offering theoretical foundations for embodied intelligent systems. Inspired by these classifications, our study proposes a novel taxonomy for visual-spatial reasoning, aiming to advance spatial reasoning research in vision-language models.
Appendix C Additional Ablation Studies
C.1 Evaluation and Format Strategy
The choice of evaluation format and answer-extraction strategy can materially influence how fully a large language model’s capabilities are expressed. Because our benchmark introduces a novel visual–spatial reasoning task, Table 5 reports the performance impact of different prompting and evaluation schemes. We consider three prompting paradigms—Direct QA, Zero-shot CoT, and Manual CoT—and three evaluation methods: (i) regex-based field extraction (RE), in which the VLM is instructed to place its answer in a fixed slot at the end of the response; (ii) JSON extraction, where the answer must appear in a dedicated answer field; and (iii) LLM-based free-form evaluation, for which we employ GPT-4.1-mini as an adjudicator.
We observe that reasoning-centric models that strongly obey formatting directives (e.g., Gemini-2.5-flash) cannot be reliably scored with regex extraction, because their compulsory chain-of-thought interferes with the required output template. For standard models, all three evaluation strategies yield comparable scores, and both CoT variants consistently outperform the Direct QA baseline by a small margin. Accordingly, in the main results table we evaluate non-reasoning models with the Manual CoT + RE setting, whereas reasoning-oriented models are assessed with Manual CoT + LLM evaluation.
Promopt Type Eval Type Avg. Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
@@ Appendix D System Prompts
We present all the system prompts used in our experiments in Figs. 7 and 8 to facilitate reproducibility. We observe that some models are sensitive to the choice of system prompt, which may stem from distributional biases in their training data. In contrast, inference-oriented models generally exhibit stronger generalization capabilities and are less reliant on carefully crafted prompts.
@@ Appendix E Additional Visualization
We present more examples of question-answer pairs from the OmniSpatial, with perspective taking, spatial interaction, dynamic reasoning, and complex logic, shown in Figs. 9, 10, 11 and 12, respectively. Our benchmark comprises a rich collection of data samples spanning diverse scenarios, resolutions, lighting conditions, and geographical regions. It includes both absolute numerical analyses and relative spatial relationships, aiming to comprehensively evaluate the spatial reasoning capabilities of vision-language models.
@@ Appendix F Limitation & Future Works
Although OmniSpatial includes some image clips with dynamic information from HOI4D [67], the complexity of the operational tasks still lags behind that of long videos [121, 14]. Moreover, while PointGraph & Spatial CoT enhances VLM’s spatial understanding through point cues, the improvement is not fundamental in nature. Spatial reasoning tasks are more like mathematics and coding tasks, require longer and more complex reasoning [22, 46].
3D information is crucial for spatial reasoning, and future work involves introducing 3D representation [23, 86, 88, 87] and perception [84, 85, 127, 30], as well as 3D VLMs [81, 24, 36, 88, 89], reasoning model [125, 22, 79] and knowledge distillation [126, 41, 23, 86]. The ultimate goal of spatial reasoning is to empower robots, and future work also involves robot execution tasks [42, 31, 89, 43, 38].

### v2

@@ 5.1 Benchmarking Spatial Reasoning
Various studies have introduced innovative benchmarking methodologies (Szymanska et al., 2024, Chen et al., 2024, Cheng et al., 2024, Cai et al., 2024, Song et al., 2024, Qi et al., 2025, Ray et al., 2024) to advance spatial reasoning evaluation. Spatial VQA (Du et al., 2024) was among the first to incorporate spatial information into vision-language models, enabling fundamental spatial relationship reasoning. SpatialBot (Cai et al., 2024) categorized spatial reasoning into various hierarchical levels, extending its applicability to robotic manipulation tasks. RoboSpatial (Song et al., 2024) proposes a large-scale template-based spatial relationship dataset, focusing on positional relationships from different perspectives. VSI-Bench (Yang et al., 2024) combined video data (Chandrasegaran et al., 2024, Liu et al., 2023c, Fu et al., 2024a, Li et al., 2024b, Fang et al., 2024b) with cognitive maps (Momennejad et al., 2023, Apostolopoulos & Groumpos, 2023) to simulate human-like spatial cognition and optimize reasoning in dynamic environments. Recently, SoFar (Qi et al., 2025) introduced the 6-DoF SpatialBench to evaluate the understanding of orientation.
While these benchmarks have made significant contributions, a unified framework encompassing a wide range of complex spatial reasoning tasks remains lacking. Inspired by prior spatial reasoning research (Xu et al., 2024b, Wang et al., 2024a, Lin et al., 2014, Nwankwo et al., 2024), we identified several limitations in existing benchmarks, such as reliance on generated images (Szymanska et al., 2024, Fu et al., 2024b, Kamath et al., 2023, Liu et al., 2023a, Rajabi & Kosecka, 2023, Shiri et al., 2024), LLM-generated templates (Linghu et al., 2024, Du et al., 2024, Rädsch et al., 2025), and domain-specific focus (Xie et al., 2025, Chow et al., 2025, Danish et al., 2024). These issues hinder their comprehensiveness and real-world applicability. To address these gaps, our study proposes a comprehensive and integrative spatial reasoning benchmark.
@@ 6 Conclusion
We introduce OmniSpatial, a benchmark for comprehensive visual–spatial reasoning. OmniSpatial distills spatial cognition into four primary categories—dynamic reasoning, complex spatial logic, spatial interaction, and perspective-taking—spanning 50 fine-grained subtasks and 8.4 K manually-curated question–answer pairs. Extensive experiments show that state-of-the-art proprietary and open-source VLMs peak at 57% accuracy—over 30 points below human performance—struggling especially with geometric reasoning and non-egocentric perspective taking. To bridge these gaps, we introduce PointGraph for structured scene-graph reasoning and SpatialCoT for viewpoint-aware chain-of-thought, both yielding consistent gains and underscoring the value of structured and multi-view reasoning.
@@ B.1 Spatial Vision-Language Models
Spatial vision-language models integrate computer vision (He et al., 2022, Oquab et al., 2024, Radford et al., 2021) and natural language processing (Bai et al., 2023, Wang et al., 2024b, Brown et al., 2020, Anil et al., 2023, Touvron et al., 2023a; b, Zhang et al., 2024b) to enhance machine understanding of spatial relationships. Recent research (Qi et al., 2024, Chen et al., 2024, Cheng et al., 2024, Cai et al., 2024, Song et al., 2024, Yuan et al., 2024, Qi et al., 2025) has increasingly focused on extending vision-language models to support spatial reasoning in dynamic and 3D environments. One of the pioneering works in this domain, SpatialVLM (Chen et al., 2024), has significantly advanced spatial reasoning by constructing an RGB-D based visual question-answering dataset. These models (Hurst et al., 2024; 2024, Reid et al., 2024, Wang et al., 2024b) effectively process multimodal data containing spatial information, serving as a bridge between visual perception and linguistic reasoning.
Further advancements (Liu et al., 2024b; a, Ramakrishnan et al., 2024, Tang et al., 2024, Yang et al., 2023, Rozanova et al., 2021, Wu et al., 2024, Stogiannidis et al., 2025) include SpatialRGPT (Cheng et al., 2024), which extends RGB-D spatial understanding by incorporating 3D scene graphs and spatial data to improve inference capabilities. Similarly, SpatialBot (Cai et al., 2024) explores hierarchical deep reasoning mechanisms to handle depth and spatial structures in complex environments. Recently, SoFar (Qi et al., 2025) proposed semantic orientation and trained an open-world orientation model to enhance the orientation understanding of VLMs, significantly improving spatial understanding and robotic operation capabilities (Cho et al., 2024, Driess et al., 2023, Mu et al., 2023, Zitkovich et al., 2023, Wake et al., 2024, Huang et al., 2024b, Nasiriany et al., 2024, Duan et al., 2024, Huang et al., 2024a, Qi et al., 2025, He et al., 2025a, Zhang et al., 2025b). However, existing research remains limited in addressing the full complexity and comprehensiveness of spatial reasoning tasks. To bridge this gap, our study aims to develop a more comprehensive benchmark that rigorously evaluates spatial reasoning capabilities.
@@ B.2 Spatial Reasoning in Psychology
In psychology, spatial reasoning refers to an individual’s ability to acquire, organize, utilize, and adapt spatial knowledge, recognized as one of the nine primary reasonings (Gardner, 2011). To systematically characterize this ability, Buckley et al. (2018)proposed a factor analysis-based framework, distinguishing between spatial visualization, spatial relations (such as mental rotation), and spatial orientation. Malanchini et al. (2020) further identified a strong correlation between spatial orientation and object manipulation skills. Additionally, Newcombe & Shipley (2014)introduced a classification system for spatial thinking, dividing spatial reasoning into two dimensions: intrinsic-extrinsic and static-dynamic. Furthermore, many studies also have contributed to the development of frameworks (Wai et al., 2009, Lee & Bednarz, 2012, Malanchini et al., 2020, Hegarty, 2010, McGarvey et al., 2018) and evaluation methods (Eliot & Smith, 1983, Uttal et al., 2024; 2013, Hegarty & Waller, 2005) for spatial reasoning. Their framework highlights the distinction between object-centric spatial properties and external reference frames, offering valuable insights for applications such as navigation and path planning. These psychological frameworks provide complementary perspectives for understanding and assessing spatial reasoning, offering theoretical foundations for embodied intelligent systems. Inspired by these classifications, our study proposes a novel taxonomy for visual-spatial reasoning, aiming to advance spatial reasoning research in vision-language models.
Promopt Type
Eval Type Avg. Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
@@
Appendix C Additional Ablation Studies
C.1 Evaluation and Format Strategy
The choice of evaluation format and answer-extraction strategy can materially influence how fully a large language model’s capabilities are expressed. Because our benchmark introduces a novel visual–spatial reasoning task, Table˜7 reports the performance impact of different prompting and evaluation schemes. We consider three prompting paradigms—Direct QA, Zero-shot CoT, and Manual CoT—and three evaluation methods: (i) regex-based field extraction (RE), in which the VLM is instructed to place its answer in a fixed slot at the end of the response; (ii) JSON extraction, where the answer must appear in a dedicated answer field; and (iii) LLM-based free-form evaluation, for which we employ GPT-4.1-mini as an adjudicator.
We observe that reasoning-centric models that strongly obey formatting directives (e.g., Gemini-2.5-flash) cannot be reliably scored with regex extraction, because their compulsory chain-of-thought interferes with the required output template. For standard models, all three evaluation strategies yield comparable scores, and both CoT variants consistently outperform the Direct QA baseline by a small margin. Accordingly, in the main results table, we evaluate non-reasoning models with the Manual CoT + RE setting, whereas reasoning-oriented models are assessed with Manual CoT + LLM evaluation.
C.2 Evaluation on the full OmniSpatial benchmark
To complement the main experiments, we further conduct evaluation on the entire 8.4K OmniSpatial test set without applying the train/test split. This setting measures model performance on all available annotated questions, thereby reducing variance due to subset sampling and providing a more holistic view of spatial reasoning ability. Table 8 reports the results across all twelve fine-grained tracks.
Overall, the relative ranking of models remains stable compared to the smaller-scale splits discussed in the main text. Gemini-2.5-Pro achieves the highest average score of 55.05, securing Rank 1 among all compared systems. ChatGPT o3 follows closely with 54.52, while Gemini-2.5-Flash achieves 51.80. Other proprietary systems, such as GPT-4.1 and GPT-4o, also maintain consistent performance. On the open-source side, Qwen-VL-2.5-32B achieves the strongest result with 43.97, outperforming other Qwen variants and aligning with the trends observed in the evaluations.
A closer per-track analysis highlights complementary strengths across models. ChatGPT o3 excels at Manipulate, Traffic Analysis, and also leads in Pattern Recognition, Ego-Centric, and Hypothetical reasoning. In contrast, Gemini-2.5-Pro dominates Locate, Geospatial Strategy, Geometric Reasoning, and Allo-Centric perspective taking. Gemini-2.5-Flash yields the highest accuracy on Motion Analysis. These complementary strengths suggest that different architectures capture distinct aspects of spatial reasoning, and point towards potential benefits of ensembling or multi-expert distillation.
Table 8: Evaluation on the full OmniSpatial benchmark. Dark green marks the best Avg. accuracy and light green marks the second-best Avg. accuracy in all the models.
Method
Avg. Rank Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
Manipulate
Motion
Analysis
Traffic
Analysis
Locate
Geospatial
Strategy
Pattern
Recognition
Geometric
Reasoning
Ego
Centric
Allo
Centric
Hypothetical
Proprietary Models
GPT-4o 44.47 5 60.00 57.97 62.38 55.24 59.09 29.90 31.01 42.06 24.34 28.05
GPT-4.1 48.75 4 61.18 66.67 66.60 69.52 60.91 30.93 35.58 47.06 27.26 33.06
o3-04-16 54.52 2 71.05 63.64 70.61 69.05 61.70 42.55 47.15 53.35 30.98 46.28
Gemini-2.0-Flash 42.41 7 56.47 56.76 55.95 63.81 57.27 15.46 30.04 43.20 24.56 28.05
Gemini-2.5-Flash 51.80 3 54.84 70.43 65.33 66.00 68.75 39.53 39.46 52.70 33.49 36.84
Gemini-2.5-Pro 55.05 1 66.67 68.34 69.10 78.85 71.15 36.17 50.17 50.65 39.57 37.89
Open-source Models
Qwen-VL-2.5-3B 39.76 8 57.65 47.34 49.07 49.52 58.18 29.90 26.97 40.62 30.47 29.88
Qwen-VL-2.5-32B 43.97 6 58.82 52.42 52.97 68.57 51.82 27.84 29.51 51.70 26.82 35.98
C.3 Details of Human Baseline & Inter-Annotator Agreement
To establish an upper bound of performance, we conducted a human evaluation on the OmniSpatial benchmark. Six human annotators (graduate students with backgrounds in vision and robotics) were recruited. Each participant was presented with a randomized subset of questions covering all four tracks and a balanced selection of fine-grained tasks. The interface was blinded: multiple-choice options were randomized and no additional context was provided. Annotators were instructed to answer independently without external resources.
Each item was labeled by three different annotators, and the final answer was obtained via majority vote. To quantify annotation reliability, we report inter-annotator agreement (IAA) using Krippendorff’s $\alpha$ for multi-annotator agreement. Table 9 summarizes the results across tracks. We also report 95% confidence intervals (bootstrap over questions) for human accuracy.
Table 9: Human baseline accuracy and inter-annotator agreement (IAA) across tracks.
Track Accuracy (%) Krippendorff’s $\alpha$
Dynamic Reasoning 95.2 $\pm$ 1.3 0.92
Spatial Interaction 93.5 $\pm$ 1.6 0.85
Complex Logic 87.9 $\pm$ 5.5 0.76
Perspective Taking 94.4 $\pm$ 2.2 0.80
Overall 92.6 $\pm$ 1.5 0.84
These results indicate that human annotators achieve $\sim$89% average accuracy, with substantial agreement across annotators ($\kappa$$\alpha\approx 0.84$). This provides a reliable estimate of the human upper bound and confirms the internal consistency of the benchmark.
@@ Appendix D System Prompts
We present all the system prompts used in our experiments in Figs.˜7 and 8 to facilitate reproducibility. We observe that some models are sensitive to the choice of system prompt, which may stem from distributional biases in their training data. In contrast, inference-oriented models generally exhibit stronger generalization capabilities and are less reliant on carefully crafted prompts.
@@ Appendix E Additional Visualization
We present more examples of question-answer pairs from the OmniSpatial, with perspective taking, spatial interaction, dynamic reasoning, and complex logic, shown in Figs.˜9, 10, 11 and 12, respectively. Our benchmark comprises a rich collection of data samples spanning diverse scenarios, resolutions, lighting conditions, and geographical regions. It includes both absolute numerical analyses and relative spatial relationships, aiming to comprehensively evaluate the spatial reasoning capabilities of vision-language models.
@@ Appendix F Limitation & Future Works
Although OmniSpatial includes some image clips with dynamic information from HOI4D (Liu et al., 2022), the complexity of the operational tasks still lags behind that of long videos (Yang et al., 2024, Chandrasegaran et al., 2024). Moreover, while PointGraph & Spatial CoT enhances VLM’s spatial understanding through point cues, the improvement is not fundamental in nature. Spatial reasoning tasks are more like mathematics and coding tasks, require longer and more complex reasoning (DeepSeek-AI et al., 2025, Jaech et al., 2024).
3D information is crucial for spatial reasoning, and future work involves introducing 3D representation (Dong et al., 2023, Qi et al., 2023a; 2024; b) and perception (Qi et al., 2017a; b, Zhang et al., 2024a, Fan et al., 2024), as well as 3D VLMs (Peng et al., 2024, Dong et al., 2024, Guo et al., 2023, Qi et al., 2024; 2025), reasoning model (Zhang et al., 2025a, DeepSeek-AI et al., 2025, OpenAI, 2025b) and knowledge distillation (Zhang et al., 2023, Hinton et al., 2015, Dong et al., 2023, Qi et al., 2023a). The ultimate goal of spatial reasoning is to empower robots, and future work also involves robot execution tasks (Huang et al., 2024a, Fang et al., 2024a, Qi et al., 2025, Huang et al., 2024b, He et al., 2025b).